# Less is More: Lean yet Powerful Vision- Language Model for Autonomous Driving

### v2

@@ 1 Introduction
Figure 2: Comparison between our framework and OpenDriveVLA-3B/7B models (Zhou et al., 2025): both methods, which adopt Qwen2.5-VL as their base model, are denoted by solid lines.
@@ 2.1 End-to-end Autonomous Driving
Conventional modular frameworks require perception modules to fully reconstruct the environment, and most end-to-end systems still follow this paradigm. However, emerging research showed that end-to-end architectures need only task-relevant perceptual features, an insight that has led to the navigation-guided implicit perception paradigm (Li & Cui, 2024), which uses driving-focused feature learning to improve inference efficiency.
@@ 3.2 Distinctions from Existing Works
The emergence of VLM-based models, such as EMMA (Hwang et al., 2024; Zhou et al., 2025), has marked a milestone in planning for autonomous driving. Although our work shares a similar goal of leveraging VLM reasoning, it diverges in several critical design philosophies that are more foundational than the specific choice of base model. These distinctions are designed to optimize the directness and efficiency of next waypoint prediction, and their key differences are as follows:
@@ 4.2 Main Results
SSR (Li & Cui, 2024) 0.19 0.36 0.62 0.39 0.25 0.64 1.33 0.74
OpenDriveVLA (Zhou et al., 2025) 0.15 0.31 0.55 0.33 0.20 0.58 1.21 0.66
@@ 5 Limitations and Future Work
  • Data Scaling and Diversity: Training on additional open-loop real-world datasets like nuPlan (H. Caesar & Tan, 2021) and Waymo Open Dataset (Sun et al., 2020) may enhance the diversity of driving styles and the model’s robustness, yet the value of incorporating unskilled driver data remains questionable.
@@ Appendix D Projection and Depth Normalization
During training, for the vision encoder, the weights of RGB channels are directly inherited from pre-trained model to preserve information, and the weights of the depth channel are Xavier-initialized (Glorot & Bengio, 2010).
@@ E.1 Basic Setups
The model was trained for approximately five days on a computing cluster equipped with $8\times$ NVIDIA A100 (80GB) GPUs.
@@ E.2 Scheduled Sampling Description and Setup
Appendix F More Visualizations

### v3

@@ 1 Introduction
Figure 2: Comparison between our framework and OpenDriveVLA-3B/7B models (Zhou et al., 2025a): both methods, which adopt Qwen2.5-VL as their base model, are denoted by solid lines.
@@ 2.1 End-to-end Autonomous Driving
Conventional modular frameworks require perception modules to fully reconstruct the environment, and most end-to-end systems still follow this paradigm. However, emerging research showed that end-to-end architectures need only task-relevant perceptual features, an insight that has led to the navigation-guided implicit perception paradigm (Li and Cui, 2024), which uses driving-focused feature learning to improve inference efficiency.
@@ 3.2 Distinctions from Existing Works
The emergence of VLM-based models, such as EMMA (Hwang et al., 2024; Zhou et al., 2025a), has marked a milestone in planning for autonomous driving. Although our work shares a similar goal of leveraging VLM reasoning, it diverges in several critical design philosophies that are more foundational than the specific choice of base model. These distinctions are designed to optimize the directness and efficiency of next waypoint prediction, and their key differences are as follows:
@@ 4.2 Main Results
SSR (Li and Cui, 2024) 0.19 0.36 0.62 0.39 0.25 0.64 1.33 0.74
AutoVLA (Zhou et al., 2025b) 0.22 0.39 0.61 0.41 0.29 0.67 1.17 0.71
OpenDriveVLA (Zhou et al., 2025a) 0.15 0.31 0.55 0.33 0.20 0.58 1.21 0.66
@@ 5 Limitations and Future Work
  • Data Scaling and Diversity: Training on additional open-loop real-world datasets like nuPlan (H. Caesar and Tan, 2021) and Waymo Open Dataset (Sun et al., 2020) may enhance the diversity of driving styles and the model’s robustness, yet the value of incorporating unskilled driver data remains questionable.
@@ Appendix D Projection and Depth Normalization
During training, for the vision encoder, the weights of RGB channels are directly inherited from pre-trained model to preserve information, and the weights of the depth channel are Xavier-initialized (Glorot and Bengio, 2010).
@@ E.1 Basic Setups
The model was trained for approximately five days over a total of 8 epochs on a computing node equipped with 8 $\times$ NVIDIA A100 (80GB) GPUs.
@@ E.2 Scheduled Sampling Description and Setup
Appendix F Case Analysis
To provide a deeper insight into our model’s capabilities beyond quantitative metrics, this section presents a qualitative analysis of its behavior in diverse scenarios. These cases illustrate that the model does not merely replicate expert trajectories but often learns and exhibits more robust, consistent, and safer driving principles, a phenomenon we interpret as emergent driving intelligence.
Case 1: Smoothness in Complex Scenarios. At this complex nighttime intersection, the model plans a fluid, graceful arc, avoiding the expert’s sharper, piecewise turn. This suggests the model learns the maneuver’s intent rather than overfitting to noisy or sub-optimal human actions.
Case 2: Conservative and Compliant Driving. At this roundabout entrance, the expert driver executes an aggressive corner-cutting maneuver. In contrast, our model generates a wider, more compliant trajectory, demonstrating that it has internalized fundamental driving rules that prioritize safety and predictability.
Case 3: Filtering Human Driver Noise. In this daytime curve, the model filters out the expert’s minor steering jitters to produce a perfectly smooth path. This highlights its ability to distill an ideal trajectory from imperfect human demonstrations, potentially leading to a more comfortable ride experience.
Case 4: Proactive and Smooth Pedestrian Evasion. While both trajectories yield to the pedestrian, the expert’s is abrupt and reactive. Our model’s response is superior: a smoother, proactive evasive maneuver initiated earlier. This demonstrates a learned defensive strategy that values predictability and comfort over reactive adjustments.
Failure Cases
Here we analyze cases where the model’s predictions diverge significantly from the ground truth. A closer look reveals these failures are rational behaviors emerging from learned driving priors, especially in scenarios where multiple valid behavioral options exist.
Case 5: Rational Choice under Uncertainty. The expert performs a left-turning straight maneuver, a decision likely based on long-term intent. Our model, lacking this context, defaults to the safest immediate action: following the lead vehicle. Its error is a rational decision under incomplete information.
Case 6: Prioritizing Stability and Consistency. The expert performs a discretionary lane change without an obvious need. The model rightly chooses to maintain its lane, a behavior that maximizes stability and predictability. This showcases a learned prior to avoid unnecessary maneuvers.
Appendix G More Visualizations